# Speaker-Smoothed kNN Speaker Adaptation for End-to-End ASR

### v1

@@ 1 Introduction
Considering the large size of E2E ASR training set, there may be utterances of similar voice characters as that of the target speaker. Thus, it is reasonable to utilize the similar utterances to be a supplement for target speaker data in the adaptation process. Motivated by this idea, we introduce a kNN retrieval techniques to ASR model. We expect the kNN classifier to find the current correctly pronounced token from its pre-built datastore at each decoding to correct the model’s erroneous output. Despite the success in natural language processing [20, 21, 22], kNN has not been researched in ASR and speaker adaptation. Therefore, we firstly explore the effectiveness of kNN representation in speaker adaptation. It was found that there may be some problems when speaker adaptation is used directly on kNN, because kNN has the fixed interpolation parameters, $T$ and $\lambda$. To solve this problem, we propose a Speaker-Smoothed kNN network, which uses x-vector and other information to dynamically adjust the interpolation ratio. [23] uses few-shot speaker data to extract similar pronunciations from training data, which is similar to our idea. However, this training method only uses utterance level information and requires speaker-related data, besides, re-training is required when the speaker is switched.
@@ 3.1 Preliminary Study
Given the limited research on the application of kNN to ASR and speaker adaptation, we initiated a preliminary study to demonstrate the potential of kNN in ASR. Utilizing Whisper, we obtained token-level representations denoted as $h_{i}$. We then sampled different speaker kNN representations across three dimensions, they are subdialect, speaker and token. The t-SNE results on $h_{i}$ embeddings from Whisper are illustrated in Figure 1. A.x samples derive from the same subdialect, where A.1 samples the same tokens with varying colors representing different speakers, while A.2 includes multiple speakers, with different color schemes indicating different tokens. B.x settings, on the other hand, sample from different subdialects, choosing up to 200 representations for four tokens or speakers. The A.x samples drawn from the same subdialect exhibit considerable overlap among speakers within the same token cluster (A.1), and a clear demarcation exists between different token clusters (A.2). These results align with our expectations and benefit our kNN retrieval. It is important to note the evident sparseness of the speaker adaptation data, as reflected in token cluster A.2 (token C), as the use of fixed temperatures and weights by kNN for interpolation increases the chance of error tokens. B.x samples, collected from different subdialects, display overlaps between clusters in B.1 and boundaries between clusters in B.2, as anticipated. However, an unfortunate overlap is found in token C of B.2. We hypothesize that this overlap may be attributed to similar pronunciations of different tokens by speakers from diverse subdialects, in addition to data sparseness.

### v3

@@ 1 Introduction
Considering the large size of E2E ASR training set, there may be utterances of similar voice characters as that of the target speaker. Thus, it is reasonable to utilize the similar utterances to be a supplement for target speaker data in the adaptation process. Motivated by this idea, we introduce a kNN retrieval techniques to ASR model which initially proposed for regulating autoregressive decoding (AD) of language model [20] and machine translation [21]. We expect this kNN classifier to find the correctly pronounced token from its pre-built datastore at each decoding to correct the model’s erroneous output. [22] modified kNN to apply to frame-level CTC decoding, but lacked further research on AD and the risk of false recall when data is sparse in kNN [23]. Therefore, we firstly explore the effectiveness of kNN token-level representation. It was found that there may be some problems when speaker adaptation is used directly on kNN, because kNN has the fixed interpolation parameters, $T$ and $\lambda$. To solve this problem, we propose a Speaker-Smoothed kNN network, which uses extra information such as x-vector to dynamically adjust the interpolation ratio. [24] uses few-shot speaker data to extract similar pronunciations from training data, which is similar to our idea. However, this training method only uses utterance level information and re-training is required when the speaker is switched.
@@ 3.1 Preliminary Study
Given the limited research on the application of kNN to ASR and speaker adaptation, we initiated a preliminary study to demonstrate the potential of kNN in ASR. Utilizing Whisper, we obtained token-level representations denoted as $h_{i}$. We then sampled different speaker kNN representations across three dimensions, they are subdialect, speaker and token. The t-SNE results on $h_{i}$ embeddings from Whisper are illustrated in Figure 1. A.x samples derive from the same subdialect, where A.1 samples the same tokens with varying colors representing different speakers, while A.2 includes multiple speakers, with different color schemes indicating different tokens. B.x settings, on the other hand, sample from different subdialects, choosing up to 200 representations for four tokens or speakers.
The A.x samples drawn from the same subdialect exhibit considerable overlap among speakers within the same token cluster (A.1), and a clear demarcation exists between different token clusters (A.2). These results align with our expectations and benefit our kNN retrieval. It is important to note the evident sparseness of the speaker adaptation data, as reflected in token cluster A.2 (token C), as the use of fixed temperatures and weights by kNN for interpolation increases the chance of error tokens.
B.x samples, collected from different subdialects, display overlaps between clusters in B.1 and boundaries between clusters in B.2, as anticipated. However, an unfortunate overlap is found in token C of B.2. We hypothesize that this overlap may be attributed to similar pronunciations of different tokens by speakers from diverse subdialects, in addition to data sparseness.